(a) Calculate the IRR, NPV, Annual Percentage Rate and Payback Period for the following projects:

PROJECT A B C D
Inicial Investment 1,000,000 2,000,000 2,000,000 1,000,000

(b) Consider the cash flow projection for the next four years. Compare the projects and determine what is the best option for an investor that wants a 10% minimum aceptable rate of return.

Years Project A Project B Project C Project D
1 300,000 400,000 400,000 1,000,000
2 400,000 200,000 200,000 1,000,000
3 500,000 100,000 100,000 -2,000,000
4 600,000 1,200,000 2,000,000
Answer #1

Cumulative PV of Cash flows (PV of current year+PV of all previous years) PV of Cash Flows(cash flows*discounting factor) CasIRR is calculted by Excel formula.

For project B and C, even the initial investment is not recovered. Therefore IRR is not available or its negative.

For A = 24.89% & D = 41.42%

NPV is simply the sum of PV of all cash flows(which also includes the initial investment as -ve figure)

It is, A = 388,771.26, B = -1,395,942.9, C = -576,326.75 & D = 598,934.5

Payback Period

= (The year in which cumulative PV crosses the initial investment-1) + Proportionate time in that year taken to recover upto initial investment

Therefore,

A = (4-1) + [(1000000-978963.19)/409808.07] = 3.05 years

B = NOT APPLICABLE (because initial investment is not recovered fully)

C = NOT APPLICABLE (because initial investment is not recovered fully)

D = (2-1) + [(1000000-909090.91)/826446.28] = 1.1 years

Annual Percentage Rate

(Assumed Simple average approach)

= (Total Return in all years/Initial Investment)/4

Therefore,

A = (1800000/1000000)/4 = 45%

B = (700000/2000000)/4 = 8.75%

C = (1900000/2000000)/4 = 23.75%

D = (2000000/1000000)/4 = 25%

From above, it can be said that project B and C should be COMPLETELY AVOIDED because of negative NPV and also nkt even the investment is recovered fully.

As far as A and C is concerned, C has more favourble NPV, IRR and Payback.

But that is because of higher payments in the initial years.

Although, C is favourable, it seems to be a HIGH RISK project, because in 3rd year outflow was as much as initial investment.

Therefore, decision should be taken considering the risks of both the project.

BEST OPTION SHOULD BE TO DIVERSIFY INTO BOTH PROJECTS, IF IT IS POSSIBLE.

(If this was helpful, kindly give positive rating to the answer. Thank You:)